OpenStudy (anonymous):

Line E has a slope of 4 FIND THE EQUATION

OpenStudy (anonymous):

y=4x. if it passes through the point 0,0

OpenStudy (anonymous):

The equation of a line is y = mx+ b (REMEMBER THIS!!!!!!!!!) y= output x = input b = y-intercept or where y=0 (x,0) m = slope so from what you have all you can determine is that y = 4x + b

OpenStudy (anonymous):

But yeah I can't stress it enough y = mx + b is important to know, as it has real world applications especially in instrumentation